clara is one third her sister's age 4 years later clara will be 3less than half her her sister 's age what was clara s age 3 years ago
Answer by lwsshak3(11628)

clara is one third her sister's age 4 years later clara will be 3less than half her her sister 's age what was clara s age 3 years ago
**
let x=clara's age 3 yrs ago
x+3=clara's present age
3(x+3)=sister's present age
4 years later:
(x+7=clara's age
(3(x+3)+4)=sister's age
3less than half of sister's age=(3(x+3)+4)/2-3
x+7=(3x+9+4)/2)-3
x+10=(3x+13)/2)
2x+20=3x+13
x=7
clara's age 3 yrs ago=7